OpenStudy (anonymous):

What are the solutions of the quadratic equation? 4x2 + 34x + 60 = 0 (1 point)

OpenStudy (anonymous):

This makes no sense.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Is "x2" supposed to be x squared?

OpenStudy (anonymous):

yeapp

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

x^2 + 11x = -24 can also be written as: x^2 + 11x +24 = 0 Now you need to ask yourself, "what two numbers add to 11 and multiply to 24?" The anser is positive 3 and positive 8. So factoring the equation yields: (x + 3)(x + 8) = 0 Set each of these individually equal to zero and solve for x: x + 3 = 0 x = -3 x + 8 = 0 x = -8 So, x = -3 and -8

OpenStudy (anonymous):

ok thank you!! let me put that in my notes

OpenStudy (anonymous):

No problem. Good luck with the rest of your work.